# Fano threefolds in positive characteristic III

### v3

@@ 5.7. Case $E_{1}-E$
The assertion (C) follows from $H_{1}\cdot H_{2}^{2}=(r_{1}-\mu_{2})^{2}L_{1}^{3}-\deg B$ (Lemma 5.11) and
@@ 5.7.1. Case $E_{1}-E_{2}$
Since each of $H_{1}$ and $H_{2}$ is nef and big, we have $H_{1}^{2}\cdot H_{2}>0$ and $H_{1}\cdot H_{2}^{2}>0$. By (B) and (C), we obtain $r_{1}>2$ and $r_{2}>1$. In particular, $r_{1}\in\{3,4\}$ and $r_{2}\in\{2,3,4\}$.
@@ 5.7.2. Cases $E_{1}-E_{3}$ and $E_{1}-E_{4}$
Since each of $H_{1}$ and $H_{2}$ is nef and big, we have $H_{1}^{2}\cdot H_{2}>0$ and $H_{1}\cdot H_{2}^{2}>0$. By (B) and (C), we obtain $r_{1}\geq 2$ and $r_{2}\geq 2$.
@@ 5.7.3. Case $E_{1}-E_{5}$
$H_{1}+H_{2}\sim-K_{X}=-f^{*}K_{X_{1}}-D_{1}=r_{1}H_{1}-D_{1},$
$H_{1}+H_{2}\sim-K_{X}=-f^{*}K_{X_{2}}-\frac{1}{2}D_{2}=\frac{r_{2}}{2}H_{2}-\frac{1}{2}D_{2}.$
@@
$H_{1}\cdot H_{2}^{2}=H_{1}\cdot((r_{1}-1)H_{1}-D_{1})^{2}=(r_{1}-1)^{2}L_{1}^{3}+H_{1}\cdot D_{1}^{2}=(r_{1}-1)^{2}L_{1}^{3}-\deg B.$
@@ 6. Primitive Fano threefolds with $\rho(X)=3$
    $(-K_{X})^{3}=14$ and $X$ is isomorphic to a prime divisor $X^{\prime}$ on $P:=\mathbb{P}(\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(-1,-1)^{\oplus 2})$ such that $\mathcal{O}_{P}(X^{\prime})\simeq\mathcal{O}_{P}(2)\otimes\pi^{*}\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(2,3)$, where $\pi:P=\mathbb{P}(\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(-1,-1)^{\oplus 2})\to\mathbb{P}^{1}\times\mathbb{P}^{1}$ denotes the natural projection.
@@
    $(-K_{X})^{3}=52$ and $X\simeq\mathbb{P}(\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(1,1))$.
@@ 6.1. Case $C-C-C$
Let $X$ be a primitive Fano threefold with $\rho(X)=3$. Assume that all the extremal rays are of type $C$. Let $F$ be a two-dimensional extremral face generated by extremal rays $R_{1}$ and $R_{2}$. Let $f_{1}:X\to Y_{1}$, $f_{2}:X\to Y_{2}$, and $\pi:X\to B$ be the contractions of $R_{1}$, $R_{2}$, and $F$, respectively. Then we have the following commutative diagram
@@
    $Y_{1}\to B$ and $Y_{2}\to B$ are the first projections.
@@
Each extremal ray $R$ is an intersection of two two-dimensional extremal faces $F$ and $F^{\prime}$. This implies that there are exactly two non-trivial contractions $X\to B_{1}$ and $X\to B_{2}$ that factor through the contraction $X\to Y$ of $R$. By $Y=\mathbb{P}^{1}\times\mathbb{P}^{1}$, these must coinside the ones induced by the projections. ∎
@@
Suppose that $X$ has at least four extrmemal rays. Then we can find two two-dimensional extremal faces $F$ and $F^{\prime}$ such that $F\cap F^{\prime}=\{0\}$. Let $\pi:X\to B$ and $\pi^{\prime}:X\to B^{\prime}$ be the contractions of $F$ and $F^{\prime}$, respectively. By Remark 6.3, we get $\dim B=\dim B^{\prime}=1$. Then there exists a curve $C$ on $X$ such that $(\pi\times\pi^{\prime})(C)$ is a point for $\pi\times\pi^{\prime}:X\to B\times B^{\prime}$. This implies that $\pi(C)$ and $\pi^{\prime}(C)$ are points, i.e., $[C]\in F\cap F^{\prime}=\{0\}$, which is absurd. ∎
@@
Since $H_{1}\cdot H_{2}$ is a fibre of $f_{12}:X\to Y_{12}=B_{1}\times B_{2}=\mathbb{P}^{1}\times\mathbb{P}^{1}$, we obtain $H_{1}\cdot H_{2}\equiv\frac{2}{\mu_{12}}\ell_{12}$, where $\mu_{12}$ denotes the length of $R_{12}$. Therefore, we get
@@
Assume $d=2$. Then $-K_{X}\sim H_{1}+H_{2}+H_{3}$ and $f:X\to Z=\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$ is a double cover (Lemma 6.6). This double cover is split (Lemma 2.4), because $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$ is $F$-split and $H^{1}(\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}}(E))=0$ for any effective divisor $E$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$. We have $(-K_{X})^{3}=\deg f\cdot\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}}(1,1,1)^{3}=12$.
Let us show that $f_{12}:X\to B_{1}\times B_{2}$ is of type $C_{1}$. We can write
@@
where ($\star$) follows from Proposition 3.16. Therefore, $b=4$. Similarly, $a=4$. Hence $f_{12}$ is of type $C_{1}$. By symmetry, $f_{ij}:X\to B_{i}\times B_{j}$ is of type $C_{1}$ for all $1\leq i<j\leq 3$. Thus (1) holds. ∎
@@ 6.2. Case $C-E$
be the contraction of $R_{i}$, where $Y_{1}=\mathbb{P}^{1}\times\mathbb{P}^{1}$. Set $D:={\operatorname{Ex}}(f_{2})$, which is a prime divisor on $X$ such that $D\simeq\mathbb{P}^{1}\times\mathbb{P}^{1}$ (Theorem 4.8). Let $\ell_{i}$ be an extremal rational curve with $R_{i}=\mathbb{R}_{\geq 0}[\ell_{i}]$. Set $\mu_{i}:=-K_{X}\cdot\ell_{i}$, which is the length of $R_{i}$.
@@
On the other hand, for $C:=f_{2}(D)$, it holds that
@@
where the first and second equalities follow from Lemma 7.2(3) and Lemma 4.7(2), respectively. We then obtain
@@ 6.2.2. Case $C_{1}-E_{1}$
We use Notation 6.9. Assume that $R_{1}$ is of type $C_{1}$. In particular, $X$ is a primitive Fano threefold with $\rho(X)=3$, $R_{1}$ is of type $C_{1}$, and $R_{2}$ is of type $E_{1}$. By Lemma 3.15, we have morphisms
${X}$${P^{\prime}:=\mathbb{P}(\mathcal{E}^{\prime})}$${Y_{1}=\mathbb{P}^{1}\times\mathbb{P}^{1}}$$\scriptstyle{\iota^{\prime}}$$\scriptstyle{\pi^{\prime}}$
where $\mathcal{E}^{\prime}:=(f_{1})_{*}\mathcal{O}_{X}(-K_{X})$ and $\iota^{\prime}$ is a closed immersion. We identify $X$ with the smooth prime divisor $\iota^{\prime}(X)$ on $P^{\prime}$.
@@
i.e., $(a-2)(b-2)=2$. Therefore, after possibly permuting the direct product factors of $Y_{1}=\mathbb{P}^{1}\times\mathbb{P}^{1}$, we obtain $(a,b)=(0,1)$ and $\mathcal{L}\simeq\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(0,1)$. Thus (1) holds. Lemma 2.4 and Lemma 8.4 imply (2) and (3), respectively.
@@
By (1) and [Tan15, Theorem 0.5], we obtain $R^{1}f_{1*}\mathcal{O}_{X}(-K_{X}-D)=0$. Thus (2) holds. This completes the proof of Step 1. ∎
@@
    $R_{1}$ is of type $C_{1}$, $R_{2}$ is of type $E_{1}$, $(-K_{X})^{3}=14$, and $X$ is isomorphic to a prime divisor $X^{\prime}$ on $P:=\mathbb{P}(\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(-1,-1)^{\oplus 2})$ such that $\mathcal{O}_{P}(X^{\prime})\simeq\mathcal{O}_{P}(2)\otimes\pi^{*}\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(2,3)$, where $\pi:P=\mathbb{P}(\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(-1,-1)^{\oplus 2})\to\mathbb{P}^{1}\times\mathbb{P}^{1}$ denotes the natural projection.
@@
See Proposition 6.11 and Proposition 6.16. ∎

### v4

@@ 5.7. Case $E_{1}-E$
The assertion (C) follows from $H_{1}\cdot H_{2}^{2}=(r_{1}-\mu_{2})^{2}L_{1}^{3}-\deg{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B_{1}}$ (Lemma 5.11) and
@@ 5.7.1. Case $E_{1}-E_{2}$
Since each of $H_{1}$ and $H_{2}$ is nef and big, we have $H_{1}^{2}\cdot H_{2}>0$ and $H_{1}\cdot H_{2}^{2}>0$. By (B) and (C), we obtain $r_{1}>2$ and $r_{2}>1$, respectively. In particular, $r_{1}\in\{3,4\}$ and $r_{2}\in\{2,3,4\}$.
@@ 5.7.2. Cases $E_{1}-E_{3}$ and $E_{1}-E_{4}$
Since each of $H_{1}$ and $H_{2}$ is nef and big, we have $H_{1}^{2}\cdot H_{2}>0$ and $H_{1}\cdot H_{2}^{2}>0$. By (B) and (C), we obtain $r_{1}\geq 2$ and $r_{2}\geq 2$, respectively.
@@
Although the following result is not used in this paper, we will need it in the classification of Fano threefolds with $\rho\geq 3$. For the definition of No. 2-xx, we refer to Definition 9.1.
Proposition 5.47.
  1. (1)
  2. (2)
Proof.
We prove (1) and (2) simultaneously. If $X$ is 2-15 (resp. 2-23), then we set $Y:=\mathbb{P}^{3}$ (resp. $Y:=Q$) and $s:=4$ (resp. $s:=3$). Then we have a blowup $f:X\to Y$ and $Y$ is a Fano threefold of index $s$. We have the contraction $g:X\to Z$ of the other extremal ray, which is of type $E_{3}$ or $E_{4}$. In particular, $T:={\operatorname{Ex}}(g)$ is a (possibly singular) quadric surface. Set $S:=f(T)$. It is enough to show that $B\subset S$ and the induced morphism
$f_{T}:T\to S$
is an isomorphism. Set $H_{f}$ (resp. $H_{g}$) to be the pullback on $X$ of the ample generator of $Y$ (resp. $Z$). It holds that $-K_{X}\sim H_{f}+H_{g}$ (Proposition 5.9). Let $r$ be the positive integer satisfying $-g^{*}K_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Z}}\sim rH_{g}$.
Step 1.
$B\subset S$.
Proof of Step 1.
$00<S\cdot C=f^{*}S\cdot C_{X}=T\cdot C_{X}<$ 0 .
This completes the proof of Step 1. ∎
Step 2.
$T\cdot\zeta=r-1$ for a one-dimensional fibre $\zeta$ of $f:X\to Y$.
Proof of Step 2.
For $E:={\operatorname{Ex}}(f)$, we have that
$-K_{X}\sim H_{f}+H_{g}\equiv\frac{-K_{X}+E}{s}+\frac{-K_{X}+T}{r},$
which implies
$-(sr-r-s)K_{X}\sim rE+sT.$
Step 3.
It holds that $r=2$ and $T\cdot\zeta=1$.
Proof of Step 3.
By [TanII, Proposition 3.22], we get
Step 4.
$f_{T}:T\to S$ is an isomorphism.
Proof of Step 4.
Let us show that $S$ is normal. Consider an exact sequence
$0\to\mathcal{O}_{X}(-T)\to\mathcal{O}_{X}\to\mathcal{O}_{T}\to 0$
By $E\cdot\zeta<0$ and $(-T-K_{X})\cdot\zeta=-1+1=0$ (Step 3), we get $R^{1}f_{*}\mathcal{O}_{X}(-T)=0$, because [Tan15, Theorem 0.5] is applicable for $A:=-T-K_{X}-\frac{1}{2}E$. Then $\mathcal{O}_{S}=f_{*}\mathcal{O}_{T}$, i.e., $S$ is normal.
Note that any curve $\zeta$ on $T$ is not contracted by $f$, as otherwise we would get a contradiction: $\zeta\in R_{1}\cap R_{2}=\{0\}$. Hence $T\neq{\operatorname{Ex}}(f)$, which implies that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{T}}:T\to S$ is birational. Since $T$ is a (possibly singular) quadric surface, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{T}}:T\to S$ is an isomorphism. This completes the proof of Step 4. ∎
Step 1 and Step 4 complete the proof of Proposition 5.47. ∎
@@ 5.7.3. Case $E_{1}-E_{5}$
$H_{1}+H_{2}\sim-K_{X}=-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{1}^{*}}K_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Y_{1}}}-D_{1}=r_{1}H_{1}-D_{1},$
$H_{1}+H_{2}\sim-K_{X}=-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{2}^{*}}K_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Y_{2}}}-\frac{1}{2}D_{2}=\frac{r_{2}}{2}H_{2}-\frac{1}{2}D_{2}.$
@@
$H_{1}\cdot H_{2}^{2}=H_{1}\cdot((r_{1}-1)H_{1}-D_{1})^{2}=(r_{1}-1)^{2}L_{1}^{3}+H_{1}\cdot D_{1}^{2}=(r_{1}-1)^{2}L_{1}^{3}-\deg{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B_{1}}.$
@@ 6. Primitive Fano threefolds with $\rho(X)=3$
    $(-K_{X})^{3}=14$ and $X$ is isomorphic to a prime divisor $X^{\prime}$ on $P:=\mathbb{P}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(-1,-1)^{\oplus 2})$ such that $\mathcal{O}_{P}(X^{\prime})\simeq\mathcal{O}_{P}(2)\otimes\pi^{*}\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(2,3)$, where $\pi:P=\mathbb{P}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(-1,-1)^{\oplus 2})\to\mathbb{P}^{1}\times\mathbb{P}^{1}$ denotes the natural projection.
@@
    $(-K_{X})^{3}=52$ and $X\simeq\mathbb{P}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(1,1))$.
@@ 6.1. Case $C-C-C$
Let $X$ be a primitive Fano threefold with $\rho(X)=3$. Assume that all the extremal rays are of type $C$. Let $F$ be a two-dimensional extremal face generated by extremal rays $R_{1}$ and $R_{2}$. Let $f_{1}:X\to Y_{1}$, $f_{2}:X\to Y_{2}$, and $\pi:X\to B$ be the contractions of $R_{1}$, $R_{2}$, and $F$, respectively. Then we have the following commutative diagram
@@
    $g_{1}:Y_{1}=B\times C_{1}\to B$ and $g_{2}:Y_{2}=B\times C_{2}\to B$ are the first projections.
@@
Each extremal ray $R$ is an intersection of two two-dimensional extremal faces $F$ and $F^{\prime}$. This implies that there are exactly two non-trivial contractions $X\to B_{1}$ and $X\to B_{2}$ that factor through the contraction $X\to Y$ of $R$. By $Y=\mathbb{P}^{1}\times\mathbb{P}^{1}$ (Theorem 4.17), these must coincide with the ones induced by the projections. ∎
@@
Suppose that $X$ has at least four extremal rays. Then we can find two two-dimensional extremal faces $F$ and $F^{\prime}$ such that $F\cap F^{\prime}=\{0\}$. Let $\pi:X\to B$ and $\pi^{\prime}:X\to B^{\prime}$ be the contractions of $F$ and $F^{\prime}$, respectively. By Remark 6.3, we get $\dim B=\dim B^{\prime}=1$. Then there exists a curve $C$ on $X$ such that $(\pi\times\pi^{\prime})(C)$ is a point for $\pi\times\pi^{\prime}:X\to B\times B^{\prime}$. This implies that $\pi(C)$ and $\pi^{\prime}(C)$ are points, i.e., $[C]\in F\cap F^{\prime}=\{0\}$, which is absurd. ∎
@@
Since $H_{1}\cdot H_{2}$ is a fibre of $f_{12}:X\to Y_{12}=B_{1}\times B_{2}=\mathbb{P}^{1}\times\mathbb{P}^{1}$, we obtain $H_{1}\cdot H_{2}\equiv\frac{2}{\mu_{12}}\ell_{12}$ (Corollary 4.5), where $\mu_{12}$ denotes the length of $R_{12}$. Therefore, we get
@@
Assume $d=2$. Then $-K_{X}\sim H_{1}+H_{2}+H_{3}$ and $f:X\to Z=\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$ is a double cover (Lemma 6.6). This double cover is split (Lemma 2.5), because $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$ is $F$-split and $H^{1}(\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}}(E))=0$ for every effective divisor $E$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$. We have $(-K_{X})^{3}=\deg f\cdot\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}}(1,1,1)^{3}=12$.
Let us show that $f_{12}:X\to B_{1}\times B_{2}$ is of type $C_{1}$ and $\Delta_{f_{12}}$ is of bidegree $(4,4)$. We can write
@@
where ($\star$) follows from Proposition 3.16. Therefore, $b=4$. Similarly, $a=4$. Hence $f_{12}$ is of type $C_{1}$ and $\Delta_{f_{12}}$ is of bidegree $(4,4)$. By symmetry, $f_{ij}:X\to B_{i}\times B_{j}$ is of type $C_{1}$ and $\Delta_{f_{ij}}$ is of bidegree $(4,4)$ for all $1\leq i<j\leq 3$. Thus (1) holds. ∎
@@ 6.2. Case $C-E$
be the contraction of $R_{i}$, where $Y_{1}=\mathbb{P}^{1}\times\mathbb{P}^{1}$ (Theorem 4.17). Set $D:={\operatorname{Ex}}(f_{2})$, which is a prime divisor on $X$ such that $D\simeq\mathbb{P}^{1}\times\mathbb{P}^{1}$ (Theorem 4.8). For each $i\in\{1,2\}$, let $\ell_{i}$ be an extremal rational curve with $R_{i}=\mathbb{R}_{\geq 0}[\ell_{i}]$. Set $\mu_{i}:=-K_{X}\cdot\ell_{i}$, which is the length of $R_{i}$.
@@
On the other hand, for $C:=f_{2}(D)(\simeq\mathbb{P}^{1})$, it holds that
@@
where the first and second equalities follow from Proposition 7.2(3) and Lemma 4.7(2), respectively. We then obtain
@@ 6.2.2. Case $C_{1}-E_{1}$
We use Notation 6.9. Assume that $R_{1}$ is of type $C_{1}$. In particular, $X$ is a primitive Fano threefold with $\rho(X)=3$, $R_{1}$ is of type $C_{1}$, and $R_{2}$ is of type $E_{1}$. By Lemma 3.15, we have a commutative diagram
${X}$${P^{\prime}:=\mathbb{P}(\mathcal{E}^{\prime})}$${Y_{1}=\mathbb{P}^{1}\times\mathbb{P}^{1}}$$\scriptstyle{\iota^{\prime}}$$\scriptstyle{f_{1}}$$\scriptstyle{\pi^{\prime}}$
where $\mathcal{E}^{\prime}:=(f_{1})_{*}\mathcal{O}_{X}(-K_{X})$, $\iota^{\prime}$ is a closed immersion, and $\pi^{\prime}$ is the natural projection. We identify $X$ with the smooth prime divisor $\iota^{\prime}(X)$ on $P^{\prime}$.
@@
i.e., $(a-2)(b-2)=2$. Therefore, after possibly permuting the direct product factors of $Y_{1}=\mathbb{P}^{1}\times\mathbb{P}^{1}$, we obtain $(a,b)=(0,1)$ and $\mathcal{L}\simeq\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(0,1)$. Thus (1) holds. Lemma 2.5(2) and Lemma 8.4 imply (2) and (3), respectively.
@@
By (1) and [Tan15, Theorem 0.5], we obtain $R^{1}(f_{1})_{*}\mathcal{O}_{X}(-K_{X}-D)=0$. Thus (2) holds. This completes the proof of Step 1. ∎
@@
    $R_{1}$ is of type $C_{1}$, $R_{2}$ is of type $E_{1}$, $(-K_{X})^{3}=14$, and $X$ is isomorphic to a prime divisor $X^{\prime}$ on $P:=\mathbb{P}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(-1,-1)^{\oplus 2})$ such that $\mathcal{O}_{P}(X^{\prime})\simeq\mathcal{O}_{P}(2)\otimes\pi^{*}\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(2,3)$, where $\pi:P=\mathbb{P}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}(-1,-1)^{\oplus 2})\to\mathbb{P}^{1}\times\mathbb{P}^{1}$ denotes the natural projection.
@@
The assertion follows from Proposition 6.11 and Proposition 6.16. ∎